# references_lipschitz_polytopes.bib

### v2

@@ Abstract.
In 2010, Vershik proposed a new combinatorial invariant of metric spaces given by a class of polytopes that arise in the theory of optimal transport and are called “Wasserstein polytopes” or “Kantorovich-Rubinstein polytopes” in the literature. Answering a question posed by Vershik, we describe the stratification of the metric cone induced by the combinatorial type of these polytopes through a hyperplane arrangement. Moreover, we study its relationships with the stratification by combinatorial type of the injective hull (i.e., the tight span) and, in particular, with certain types of metrics arising in phylogenetic analysis. We also compute enumerative invariants in the case of metrics on up to six points.
@@ 1.1. The main results
    We introduce a Wasserstein arrangement whose cells correspond to the combinatorial types of KRW polytopes. This arrangement builds upon the Gordon–Petrov description of the faces of KRW polytopes and could be of independent interest. Moreover, via the above connection to alcoved polytopes, intersecting this arrangement with the metric cone naturally yields the type-fan of symmetric alcoved polytopes.
@@
    Using the computer algebra system OSCAR [OSCAR-book] together with the Julia package CountingChambers.jl [BEK23] and the software TOPCOM [Rambau] we obtain an enumeration of the combinatorial types of KRW polytopes of generic metrics on $n=4,5,6$ points. This is displayed in Table 1, where the number of combinatorial types of generic KRW polytopes – i.e., the number of chambers of the Wasserstein arrangement – is shown in the column titled “Labeled”. The column “Unlabeled” shows the number of combinatorial types up to combinatorial isomorphisms (where we are allowed to permute the vertices to obtain a combinatorial isomorphism), which is the number of chambers of the Wasserstein arrangement modulo the natural action of the symmetric group $S_{n}$ via coordinate permutation.
@@
    $n$ Unlabeled Labeled
    3 1 1
    4 1 6
    5 12 882
    6 25,224 6,677,863,200
    Table 1. Combinatorial types of KRW polytopes of generic $n$-metrics.
@@ Acknowledgments
We thank Nick Early and Benjamin Schröter for fruitful discussions, as well as Laura Casabella and Michael Joswig for helpful comments on earlier versions of this manuscript. We are grateful to Jörg Rambau for providing us with relevant triangulations of the root polytope computed using TOPCOM.
LK and LM are supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB-TRR 358/1 2023 – 491392403 and SPP 2458 – 539866293.
@@ 2.1. The metric cone
$x_{\{i,j\}}>0,\quad x_{\{i,j\}}+x_{\{j,k\}}\geq x_{\{i,k\}},\textrm{ for all pairwise distinct }i,j,k\in[n]$
. The pseudometric cone, denoted by $\overline{\mathbf{M}}_{n}$, is the (topological) closure of $\mathbf{M}_{n}$.
@@ 2.2. Admissible graphs and strict metrics
Let $\rho$ be strict and consider an instance of Theorem 2.5.(2) with $x_{i}=y_{j}$ for some $i$ and $j$. Then (2.1) applied to $(x_{j},y_{j})$, $(x_{i},y_{i})=(y_{j},y_{i})$ gives
$\rho_{xy}+\rho_{yz}\leq\rho_{yy}+\rho_{xz}=\rho_{xz}$
and thus, $\rho_{xy}+\rho_{yz}=\rho_{xz}$ which contradicts the assumption on $\rho$ being strict. ∎
@@ 2.3. Polyhedra and their combinatorial type
We refer to [ZieglerLoP] for basics and terminology about polyhedra and only recall some fundamental definitions. A polyhedron is any subset of Euclidean space that is obtained intersecting a finite number of closed halfspaces. A polytope is a bounded polyhedron – equivalently, the convex hull of a finite set of points. Every polyhedron is topologically closed. The affine hull of a polytope $P$ is the smallest affine subspace containing $P$, and the relative interior of $P$ is the interior of $P$ as a subset of its affine hull. If $P$ is any convex set in Euclidean space, we call face of $P$ any intersection of $P$ with the boundary of a closed halfspace containing $P$.
@@
Example 2.11.
If $\rho$ is an $n$-metric, the polytope $KRW(\rho)$ together with all its faces defines a polyhedral complex, whose vertices are naturally labeled by ordered pairs of distinct elements of $[n]$ (i.e., if $p_{ij}$ is a vertex of $KRW(\rho)$, it is labeled by $(i,j)$). By Section 2.2, the KRW polytopes of two metrics $\rho_{1}$, $\rho_{2}$ have the same labeled combinatorial type if and only if the set of admissible labeled graphs for $\rho_{1}$ concides with the set of admissible labeled graphs for $\rho_{2}$, where the labeling of the graphs is induced by the labeling of the edges $(i,j)$, which are in one-to-one correspondence with the vertices $p_{ij}$.
We will call cone any subset $\mathbf{C}\subseteq R^{d}$ that contains every linear combination with positive coefficients of any of its finite subsets. Note that we thus slightly deviate from the standard terminology where zero coefficients are allowed, as in [ZieglerLoP, §1.1] in order to preserve the term ”metric cone” for $\mathbf{M}_{n}$. We call polyhedral cone any polyhedron containing 00 that is also a cone (this is equivalent to the standard notion).
Following standard terminology, we call polyhedral fan any polyhedral complex all of whose members are polyhedral cones, see [ZieglerLoP, §7.1]. We will call fan any set of (not necessarily polyhedral) cones that contains all faces of each of its members, and such the intersection of any two members is a face of both: again, this is not standard but allows us, e.g., to speak of the ”Wasserstein fan” $\mathscr{F}_{n}$ in Definition 3.3.
@@ 3. The Wasserstein arrangement and the Wasserstein fan
Let $n\geq 2$ be an integer and consider two $n$-metrics $\rho^{(1)},\rho^{(2)}$. The following two statements are equivalent:
@@
    The metrics $\rho^{(1)},\rho^{(2)}$ are in the open part of one cone $\sigma$ of the Wasserstein fan $\mathscr{F}_{n}$.
@@
    The combinatorial structures of the polytopes $KRW(\rho^{(1)})$ and $KRW(\rho^{(2)})$ are the same, i.e., both polytopes have the same set of admissible directed graphs.
@@
Following Theorem 2.5, the sign pattern of a metric on an open cone in the fan $\mathscr{F}_{n}$, i.e., describing whether the metric lies on, “under” or “above” a hyperplane, specifies the collection of admissible graphs of the metric. As the latter is the combinatorial structure of the KRW polytope of the metric, the statement follows. ∎
@@ 4. Regular triangulations and the secondary fan
The next proposition establishes a relationship between regular subdivisions of $\mathcal{R}_{n}^{0}$ and KRW polytopes. A similar statement also appeared in the second arXiv version of the preprint [JS19].
@@
$\mathscr{F}_{n}=\pi\big{(}\Sigma(\mathcal{R}_{n}^{0})\cap Q\big{)}\cap\mathbf{M}_{n},$
@@
Let $\rho$ be an $n$-metric in $\mathbf{M}_{n}$ and let $F_{\rho}$ be the unique open face of the Wasserstein fan $\mathscr{F}_{n}$ that contains $\rho$. Furthermore, let $\rho^{\prime}$ be another $n$-metric. By construction of the regular subdivisions (Proposition 4.3.(1)) and by symmetry of $\rho$ and $\rho^{\prime}$, surely $\mathcal{R}_{n}^{0}(\rho)$ and $\mathcal{R}_{n}^{0}(\rho^{\prime})$ lie in $Q$.
Now, by construction of $\mathscr{F}_{n}$ and Theorem 2.5, the polytopes $KRW(\rho)$ and $KRW(\rho^{\prime})$ determine the same collection of admissible graphs and therefore have the same combinatorial structure if and only if $\rho^{\prime}\in F_{\rho}$. By Proposition 4.3, this means that $\mathcal{R}_{n}^{0}(\rho)$ and $\mathcal{R}_{n}^{0}(\rho^{\prime})$ define the same regular subdivision of $\mathcal{R}_{n}^{0}$ if and only if $\rho^{\prime}\in F_{\rho}$. Hence $\rho$ and $\rho^{\prime}$ lie in the same open cone of $\Sigma(\mathcal{R}_{n}^{0})\cap Q$ if and only if $\rho^{\prime}\in F_{\rho}$ which implies the claim. ∎
@@ 5. Relationship to tight spans and injective hulls
We will speak of the labeled combinatorial type of tight spans according to the terminology of Example 2.11 with respect to the following labeling of vertices. Some vertices of $E(\rho)$ have a natural labeling by elements of $[n]$ discussed in Remark 5.4. When passing to $E^{\sharp}(\rho)$ via the characterization given in Remark 5.6, for each $k$ such that $h^{(k)}$ is pruned, there is a unique vertex of $E^{\sharp}(\rho)$, called $v^{(k)}$ in Theorem A.3, that is incident to the pruned edge. Note that if $\rho$ is a metric, for different $k$ all $h^{(k)}$ are distinct, but some $v^{(k)}$ may concide. Thus we can label vertices of $E^{\sharp}(\rho)$ either by the label they inherit from $E(\rho)$ or by the set of labels of adjacent pruned vertices in $E(\rho)$.
@@
The fan $\mathscr{F}_{n}$ is neither a coarsening nor a refinement of the subdivision $\mathscr{S}_{n}$ of the metric cone induced by secondary fan of $\Delta_{2,n}$ for $n\geq 5$. Similarly, the fan ¯$\overline{\mathscr{F}}_{n}$ is neither a coarsening nor a refinement of the metric fan $\operatorname{MF}(n)$ for $n\geq 5$.
@@
The fan $\mathscr{F}_{4}$ is a refinement of the subdivision $\mathscr{S}_{4}$ of the metric cone induced by the secondary fan of $\Delta_{2,4}$. See Figure 8 and Section 8.1.
@@
For every $n\geq 4$ there are strict metrics for which the injective hulls have the same labeled combinatorial type, but whose KRW-polytopes have different unlabeled combinatorial type. For $n\geq 5$ there are strict metrics for which the KRW-polytopes have the same labeled combinatorial type, but the injective hulls have different unlabeled combinatorial type.
@@
denote the image of $h^{(n)}\in E(\rho)$ in $E(\sigma)$, and let [^1] Sum of subsets of Euclidean space is intended element-wise, i.e., as Minkowski sum.
@@
Let $\rho_{1}$, $\rho_{2}$ be the two $5$-metrics of Example 6.13. The isomorphism between the polytopes $KRW(\rho_{1})$ and $KRW(\rho_{2})$ induces an isomorphism between $KRW(\rho_{1}^{(k)})$ and $KRW(\rho_{2}^{(k)})$. On the other hand, any isomorphism between the tight spans of $\rho_{1}^{(k)}$ and $\rho_{2}^{(k)}$ must map the $t^{(n,k)}$-part to itself (since the tight spans of $\rho_{1}$ and $\rho_{2}$ do not contain any path of length more than $1$ consisting of vertices with degree $\leq 2$, see Example 6.13), and thus would induce an isomorphism between the tight spans of $\rho_{1}$ and $\rho_{2}$. The latter cannot exist, and so we show that $\mathscr{F}_{n}$ is not a refinement of the secondary fan of $\Delta_{2,n}$ for all $n\geq 5$.

### v3

@@ Abstract.
In 2010, Vershik proposed a new combinatorial invariant of metric spaces given by a class of polytopes that arise in the theory of optimal transport and are called “Wasserstein polytopes” or “Kantorovich-Rubinstein polytopes” in the literature. Answering a question posed by Vershik, we describe the stratification of the metric cone induced by the combinatorial type of these polytopes via an associated hyperplane arrangement. Moreover, we study its relationships with the stratification by combinatorial type of the injective hull (i.e., the tight span) and, in particular, with certain types of metrics arising in phylogenetic analysis. We also compute enumerative invariants in the case of metrics on up to six points.
@@ 1.1. The main results
    The subdivision of the metric cone by combinatorial types of KRW polytopes is a fan whose cones are (interiors of) polyhedral cones: we call this the type fan of KRW polytopes. We introduce a Wasserstein arrangement which refines the type fan of KRW polytopes. This arrangement builds upon the Gordon–Petrov description of the faces of KRW polytopes and could be of independent interest. Moreover, via the above connection to alcoved polytopes, intersecting this arrangement with the metric cone naturally yields a refinement of the type fan of symmetric alcoved polytopes.
@@
    Using the computer algebra system OSCAR [OSCAR-book] together with the Julia package CountingChambers.jl [BEK23] and the software TOPCOM [Rambau] we obtain an enumeration of the combinatorial types of KRW polytopes of generic metrics on $n=4,5,6$ points. This is displayed in Table 1, where the number of combinatorial types of generic KRW polytopes – i.e., the number of chambers of the Wasserstein arrangement – is shown in the column titled “Labeled”. The column “Unlabeled” shows the number of combinatorial types up to combinatorial isomorphisms (where we are allowed to permute the vertices to obtain a combinatorial isomorphism). The last column of this table shows the number of chambers of the Wasserstein arrangement which agrees with the number of combinatorial type of generic KRW polytopes up to $n\leq 5$. This stems from the fact that for $n\geq 6$ the Wasserstein arrangement is a strict refinement of the type fan of KRW polytopes, see Proposition 3.13.
@@
    $n$ Unlabeled Labeled # Chambers
    3 1 1 1
    4 1 6 6
    5 12 882 882
    6 25,224 17,695,320 6,677,863,200
    Table 1. Combinatorial types of KRW polytopes of generic $n$-metrics and chambers of the Wasserstein arrangement.
@@ Acknowledgments
We thank Nick Early and Benjamin Schröter for fruitful discussions, as well as Laura Casabella and Michael Joswig for helpful comments on the manuscript. We are grateful to Jörg Rambau for providing us with relevant triangulations of the root polytope computed using TOPCOM. We would like to thank Alexander Postnikov for pointing out a mistake in an earlier version of the article.
LK and LM are funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 491392403 – TRR 358 and SPP 2458 – 539866293.
@@ 2.1. The metric cone
$x_{\{i,j\}}>0,\quad x_{\{i,j\}}+x_{\{j,k\}}\geq x_{\{i,k\}},\textrm{ for all pairwise distinct }i,j,k\in[n].$
The pseudometric cone, denoted by $\overline{\mathbf{M}}_{n}$, is the (topological) closure of $\mathbf{M}_{n}$.
@@ 2.2. Admissible graphs and strict metrics
Let $\rho$ be strict and consider an instance of Theorem 2.5.(2) with $x_{i}=y_{j}$ for some distinct $i$ and $j$. Then (2.1) applied to $(x_{j},y_{j})$, $(x_{i},y_{i})=(y_{j},y_{i})$ gives
$\rho_{x_{j}y_{j}}+\rho_{x_{i}y_{i}}\leq\rho_{x_{j}y_{i}}+\rho_{y_{j}x_{i}}=\rho_{x_{j}y_{i}}$
and thus, $\rho_{x_{j}y_{j}}+\rho_{y_{j}y_{i}}=\rho_{x_{j}y_{i}}$ which contradicts the assumption on $\rho$ being strict. ∎
@@ 2.3. Polyhedra and their combinatorial type
We refer to [ZieglerLoP] for basics and terminology about polyhedra and only recall some fundamental definitions. A polyhedron is any subset of Euclidean space that is obtained by intersecting a finite number of closed halfspaces. A polytope is a bounded polyhedron – equivalently, the convex hull of a finite set of points. Every polyhedron is topologically closed. The affine hull of a polytope $P$ is the smallest affine subspace containing $P$, and the relative interior of $P$ is the interior of $P$ as a subset of its affine hull. If $P$ is any convex set in Euclidean space, we call face of $P$ any intersection of $P$ with the boundary of a closed halfspace containing $P$.
@@
Remark 2.11.
If $\rho$ is an $n$-metric, the polytope $KRW(\rho)$ together with all its faces defines a polyhedral complex, whose vertices are naturally labeled by ordered pairs of distinct elements of $[n]$ (i.e., if $p_{ij}$ is a vertex of $KRW(\rho)$, it is labeled by $(i,j)$). Note that the ”combinatorial structure” in the sense of Gordon and Petrov, see Section 2.2, is equivalent to the labeled combinatorial type. Indeed, by Theorem 2.5 the KRW polytopes of two metrics $\rho_{1}$, $\rho_{2}$ have the same labeled combinatorial type if and only if the set of admissible labeled graphs for $\rho_{1}$ coincides with the set of admissible labeled graphs for $\rho_{2}$, where the labeling of the graphs is induced by the labeling of the edges $(i,j)$, which are in one-to-one correspondence with the vertices $p_{ij}$.
We will call cone any subset $\mathbf{C}\subseteq\mathbb{R}^{d}$ that contains every linear combination with positive coefficients of any of its finite subsets. Note that we thus slightly deviate from the standard terminology where zero coefficients are allowed, as in [ZieglerLoP, §1.1] in order to preserve the term ”metric cone” for $\mathbf{M}_{n}$. We call polyhedral cone any polyhedron containing 00 that is also a cone (this is equivalent to the standard notion).
Following standard terminology, we call polyhedral fan any polyhedral complex all of whose members are polyhedral cones, see [ZieglerLoP, §7.1]. We will call fan any set of (not necessarily polyhedral) cones that contains all faces of each of its members, and such the intersection of any two members is a face of both: again, this is not standard but allows us, e.g., to speak of the ”Wasserstein fan” $\mathscr{F}_{n}$ in Definition 3.3
Definition 2.12.
Let $n\geq 2$ be an integer. We write $\mathscr{T}_{n}$ for the type fan of KRW polytopes, i.e., the fan supported on the metric cone $\mathbf{M}_{n}$ such that two metrics $\rho^{(1)},\rho^{(2)}$ are in the open part of the same cone of $\mathscr{T}_{n}$ if and only if the the KRW polytopes $KRW(\rho^{(1)})$ and $KRW(\rho^{(2)})$ have the same labeled combinatorial type (equivalently, the same set of admissible directed graphs).
Remark 2.13.
Note that $\mathscr{T}_{n}$ is indeed a fan, and the relatively open part of its cones is the relative interior of a polyhedral cone. In fact, according to Theorem 2.5 the region within the metric cone consisting of the metrics whose KRW polytopes have a fixed combinatorial type is cut out by a collection of linear inequalities.
@@ 3. The Wasserstein arrangement and the Wasserstein fan
Let $n\geq 2$ be an integer and consider two $n$-metrics $\rho^{(1)},\rho^{(2)}$. If the metrics $\rho^{(1)},\rho^{(2)}$ are in the open part of one cone of the Wasserstein fan $\mathscr{F}_{n}$ then the labeled combinatorial type of the polytopes $KRW(\rho^{(1)})$ and $KRW(\rho^{(2)})$ are the same, i.e., $\rho^{(1)},\rho^{(2)}$ determine the same set of admissible directed graphs.
Hence the Wasserstein fan $\mathscr{F}_{n}$ refines the type fan of KRW polytopes $\mathscr{T}_{n}$.
Proof.
Following Theorem 2.5, the sign pattern of a metric in an open cone in the fan $\mathscr{F}_{n}$, i.e., describing whether the metric lies on, “under” or “above” a hyperplane, specifies the collection of admissible graphs of the metric. As noted in Remark 2.11, this data determines the labeled combinatorial type. ∎
We next study the relationship between these two fans further.
Proposition 3.13.
Let $n\geq 2$ and let $\mathcal{W}_{n}^{4}$ denote the subset of $\mathcal{W}_{n}$ consisting of all hyperplanes of the form $H_{(i,j),(k,l)}$ for distinct $i,j,k,l\in[n]$.
@@
    For every $H\in\mathcal{W}_{n}^{4}$, the set $H\cap\mathbf{M}_{n}$ is a union of cones of the type fan $\mathscr{T}_{n}$.
@@
    The union of all $H\cap\mathbf{M}_{n}$, where $H$ runs over all elements of $\mathcal{W}_{n}^{4}$, supports a subfan of the Wasserstein fan $\mathscr{F}_{n}$. This subfan is a refinement of a subfan of the type fan $\mathscr{T}_{n}$.
  3. (3)
    For $n\leq 5$, the Wasserstein fan $\mathscr{F}_{n}$ agrees with the type fan $\mathscr{T}_{n}$.
  4. (4)
    For $n\geq 6$, the Wasserstein fan $\mathscr{F}_{n}$ is a strict refinement of the type fan $\mathscr{T}_{n}$.
@@
Let $\rho$ be an $n$-metric, fix distinct $i,j,k,l\in[n]$ and consider the directed graphs $G^{+}$, $G^{-}$ on the vertex set $[n]$ with edge sets, respectively, $\{(i,j),(k,l)\}$ and $\{(i,l),(k,j)\}$. By Theorem 2.5, $G^{+}$ is admissible if and only if $\rho_{ij}+\rho_{kl}\leq\rho_{ik}+\rho_{jl}$ (i.e., $\rho\in H_{(i,j),(k,l)}^{+}$), and $G^{-}$ is admissible if and only if $\rho_{ij}+\rho_{kl}\geq\rho_{ik}+\rho_{jl}$ (i.e., $\rho\in H_{(i,j),(k,l)}^{+}$). Moreover, for $\epsilon=+,-$ the graph $G^{\epsilon}$ is admissible for $\rho$ if and only if it is admissible for every metric whose KRW polytope has the same labeled combinatorial type as $KRW(\rho)$ (see, e.g., Remark 2.11). Thus, the open cone of $\mathscr{T}_{n}$ containing $\rho$ is either contained in $H_{(i,j),(k,l)}$ (in case both $G^{+}$ and $G^{-}$ are admissible) or disjoint from it. This proves (1).
For item (2) write $W^{(4)}$ for the union of all $H\cap\mathbf{M}_{n}$ with $H\in\mathcal{W}_{n}^{4}$. Then $W^{(4)}$ clearly supports a subfan of $\mathscr{F}_{n}$ since $\mathcal{W}_{n}^{4}\subseteq\mathcal{W}_{n}$. Moreover, part (1) and the fact that $W^{(4)}$ is closed in $\mathbf{M}_{n}$ imply that $W^{(4)}$ also supports a subfan of $\mathscr{T}_{n}$. Noticing that every cone of $\mathscr{T}_{n}$ is a union of cones of $\mathscr{F}_{n}$ completes the proof of part (2).
For $n\leq 6$ the statements (3) and (4) follow from Proposition 3.12 and Table 1. In order to complete the proof of (4), we let $n\geq 6$ and consider the triple of edges $T=\{(1,2),(3,4),(5,6)\}$. According to Theorem 2.5, the directed graph defined by $T$ is admissible if and only if the following two inequalities hold:
$X:=\rho_{12}+\rho_{34}+\rho_{56}\leq\rho_{16}+\rho_{45}+\rho_{23}:=Y,$
$X:=\rho_{12}+\rho_{34}+\rho_{56}\leq\rho_{14}+\rho_{36}+\rho_{25}:=Z.$
The Wasserstein arrangement contains the three hyperplanes $X=Y$, $X=Z$, and $Y=Z$. These subdivide $\mathbf{M}_{n}$ in $6$ chambers, in two of which the collection $T$ is admissible. On the other hand, the type fan is subdivided by these conditions into $3$ cells, according to which of the three terms $X$, $Y$, and $Z$ is minimal (this is thus a tropical hyperplane). The cell in which $X$ is minimal is the union of the two above chambers and the region in which $T$ is admissible. Thus, already for these subfans we see that the Wasserstein subfan is a strict refinement of the type subfan as every cell of the latter is subdivided into two pieces by the former. Choosing a generic point on one of these separating wall of the Wasserstein fan and moving infinitesimally in the direction above and below this wall yields two points $\mathbf{M}_{n}$ such that both points are in the open part of one cone of the type fan $\mathscr{T}_{n}$ but these two points are in different chambers of the Wasserstein arrangement and thus also in different open cells of the Wasserstein fan $\mathscr{F}_{n}$. ∎
@@ 4. Regular triangulations and the secondary fan
The next proposition establishes a relationship between regular subdivisions of $\mathcal{R}_{n}^{0}$ and KRW polytopes. A similar statement also appeared in the second arXiv version of the preprint [JS19] and more recently in [AF24, Theorem 3.8].
@@
$\mathscr{T}_{n}=\pi\big{(}\Sigma(\mathcal{R}_{n}^{0})\cap Q\big{)}\cap\mathbf{M}_{n},$
@@
Let $\rho$ be an $n$-metric in $\mathbf{M}_{n}$ and let $F_{\rho}$ be the unique open face of the type fan $\mathscr{T}_{n}$ that contains $\rho$. Furthermore, let $\rho^{\prime}$ be another $n$-metric. By construction of the regular subdivisions (Proposition 4.3.(1)) and by symmetry of $\rho$ and $\rho^{\prime}$, surely $\mathcal{R}_{n}^{0}(\rho)$ and $\mathcal{R}_{n}^{0}(\rho^{\prime})$ lie in $Q$.
Now, by construction of $\mathscr{T}_{n}$ the polytopes $KRW(\rho)$ and $KRW(\rho^{\prime})$ determine the same combinatorial structure if and only if $\rho^{\prime}\in F_{\rho}$. By Proposition 4.3, this means that $\mathcal{R}_{n}^{0}(\rho)$ and $\mathcal{R}_{n}^{0}(\rho^{\prime})$ define the same regular subdivision of $\mathcal{R}_{n}^{0}$ if and only if $\rho^{\prime}\in F_{\rho}$. Hence $\rho$ and $\rho^{\prime}$ lie in the same open cone of $\Sigma(\mathcal{R}_{n}^{0})\cap Q$ if and only if $\rho^{\prime}\in F_{\rho}$ which implies the claim. ∎
@@ 5. Relationship to tight spans and injective hulls
We will speak of the labeled combinatorial type of tight spans according to the terminology of Remark 2.11 with respect to the following labeling of vertices. Some vertices of $E(\rho)$ have a natural labeling by elements of $[n]$ discussed in Remark 5.4. When passing to $E^{\sharp}(\rho)$ via the characterization given in Remark 5.6, for each $k$ such that $h^{(k)}$ is pruned, there is a unique vertex of $E^{\sharp}(\rho)$, called $v^{(k)}$ in Theorem A.3, that is incident to the pruned edge. Note that if $\rho$ is a metric, for different $k$ all $h^{(k)}$ are distinct, but some $v^{(k)}$ may concide. Thus we can label vertices of $E^{\sharp}(\rho)$ either by the label they inherit from $E(\rho)$ or by the set of labels of adjacent pruned vertices in $E(\rho)$.
@@
For $n\geq 5$, the fans $\mathscr{F}_{n}$ and $\mathscr{T}_{n}$ are neither coarsenings nor refinements of the subdivision $\mathscr{S}_{n}$ of the metric cone induced by the secondary fan of $\Delta_{2,n}$. Similarly, for $n\geq 5$ the fan ¯$\overline{\mathscr{F}}_{n}$ is neither a coarsening nor a refinement of the metric fan $\operatorname{MF}(n)$.
@@
The fan $\mathscr{F}_{4}=\mathscr{T}_{4}$ is a refinement of the subdivision $\mathscr{S}_{4}$ of the metric cone induced by the secondary fan of $\Delta_{2,4}$. See Figure 8 and Section 8.1.
@@
For every $n\geq 4$ there is a pair of strict metrics whose injective hulls have the same labeled combinatorial type, but whose KRW-polytopes have different unlabeled combinatorial type. For $n\geq 5$ there is a pair of strict metrics for which the KRW-polytopes lie in the same open cell of the Wasserstein fan $\mathscr{F}_{n}$ (and hence they have the same labeled combinatorial type), but whose injective hulls have different unlabeled combinatorial type.
@@
denote the image of $h^{(n)}\in E(\rho)$ in $E(\sigma)$, and let [^1] Sum of subsets of Euclidean space is intended elementwise, i.e., as Minkowski sum.
@@
Let $\rho_{1}$, $\rho_{2}$ be the two $5$-metrics of Example 6.13 which lie in the same cone of the Wasserstein fan $\mathscr{F}_{5}$. The isomorphism between the polytopes $KRW(\rho_{1})$ and $KRW(\rho_{2})$ induces an isomorphism between $KRW(\rho_{1}^{(k)})$ and $KRW(\rho_{2}^{(k)})$. On the other hand, any isomorphism between the tight spans of $\rho_{1}^{(k)}$ and $\rho_{2}^{(k)}$ must map the $t^{(n,k)}$-part to itself (since the tight spans of $\rho_{1}$ and $\rho_{2}$ do not contain any path of length more than $1$ consisting of vertices with degree $\leq 2$, see Example 6.13), and thus would induce an isomorphism between the tight spans of $\rho_{1}$ and $\rho_{2}$. The latter cannot exist, and so we show that $\mathscr{F}_{n}$ is not a refinement of the secondary fan of $\Delta_{2,n}$ for all $n\geq 5$.